# Geometric Integration by Parts and Sobolev Spaces on Vector Bundles: A Unified Global Approach

### v1

@@ Abstract
This article develops a unified framework for the theory of Sobolev spaces on vector bundles over riemannian manifolds. The analytical core of our approach is a rigorous higher-order geometric integration by parts formula, which characterizes the formal adjoint of the covariant derivative. This identity is established for arbitrary manifolds, requiring no assumptions on completeness or compactness. While these results are fundamental to global analysis, explicit and direct proofs are often elusive in the literature or rely on overly sophisticated machinery that overshadows the underlying geometry. To bridge this gap, we establish sharp local-to-global norm equivalence estimates and provide streamlined, self-contained proofs for the Meyers-Serrin theorem on general manifolds, as well as the Sobolev embedding and Rellich-Kondrashov theorems for the compact case. By prioritizing intrinsic global arguments over ad hoc coordinate patching, this work provides a modern and accessible foundation for the study of Sobolev spaces on bundles.
Geometric Integration by Parts and Sobolev Spaces on Vector Bundles: A Unified Global Approach
Velázquez-Mendoza Carlos Daniel
Department of Mathematics. Facultad de Ciencias, UNAM. Circuito Exterior s/n. México City. 04510. México.
danielc21@ciencias.unam.mx
Sandoval-Romero María de los Ángeles (Corresponding Author)
Department of Mathematics. Facultad de Ciencias, UNAM. Circuito Exterior s/n. México City. 04510. México.
selegna@ciencias.unam.mx
@@
Keywords: Sobolev spaces, Vector bundles, Geometric integration by parts, Embedding Theorems, Rellich-Kondrashov Theorem.
1.  Introduction
@@
The analysis of Sobolev spaces on vector bundles is a foundational pillar of modern geometric analysis. A central issue in this theory is ensuring the equivalence between the two standard definitions of these spaces: the completion of the space of smooth sections with respect to the Sobolev norm, denoted by $H^{m,p}(E)$, and the space of sections possessing weak covariant derivatives in $L^{p}$, denoted by $W^{m,p}(E)$.
In Riemannian manifolds, the treatment of this equivalence often varies significantly across the literature. Standard texts such as [1] and [4] typically define Sobolev spaces directly as completions of smooth sections, largely bypassing the intrinsic difficulties of weak derivatives. In contrast, approaches that prioritize intrinsic geometry, such as [7], define them via distributional derivatives. Establishing the identity $H^{m,p}(E)=W^{m,p}(E)$—the Meyers-Serrin theorem—is therefore essential to reconcile these perspectives. While this equivalence is well-documented for scalar functions or under specific geometric assumptions (see [2]), the treatment for general vector bundles requires rigorous handling of the fiber metric and the connection.
Existing proofs of the fundamental structural theorems—density, embedding, and compactness—often suffer from a dichotomy between generality and geometric clarity. For instance, the density of smooth sections is established in [3] for arbitrary vector bundles, but the framework relies on the theory of general linear partial differential operators and requires local elliptic regularity. Conversely, regarding the Sobolev embedding and Rellich-Kondrashov theorems, the seminal work of Palais [8] remains the most general reference for compact manifolds. Despite its importance, the proofs in [8] involve a complex reduction to the $n$-dimensional torus that can obscure the intrinsic geometry of the manifold.
The purpose of this paper is to provide a unified, constructive, and strictly geometric framework for the theory of Sobolev spaces on vector bundles. We avoid reliance on elliptic regularity or indirect embeddings into tori. Instead, we develop the theory from the intrinsic definition of the weak covariant derivative, utilizing a rigorous integration by parts formula as the analytical core.
Our approach yields three main contributions: First, in Section 3, we derive a higher-order integration by parts formula for the covariant derivative acting on sections of vector bundles (Theorem 1). Beyond its global significance, this identity provides a constructive path to characterize the formal adjoint of the covariant derivative. By working in arbitrary local frames, we expose the precise dependence of the adjoint on the structural properties of the connection and the fiber metrics, resolving the combinatorial complexity that arises at higher orders. Second, in Section 5, we present a self-contained proof of the Meyers–Serrin theorem valid for general Riemannian manifolds. Unlike previous approaches, our proof rests solely on the geometric integration by parts formula and standard convolution arguments within local charts, bridged by sharp local-to-global norm equivalence estimates. Finally, for the case of compact manifolds, in Section 6, we utilize this framework to provide streamlined proofs of the continuous Sobolev embedding theorem and the Rellich–Kondrashov compactness theorem. By prioritizing intrinsic global arguments over ad hoc coordinate patching, this work offers a rigorous foundation for global analysis compatible with modern geometric applications.
2.  Notation and Preliminaries
@@ 2.2 Sobolev Spaces
We begin with the precise definitions for the Sobolev spaces used in this work.
@@
Recall that for the tensor field $F^{(a)}$, the connection components are given by (see, e.g., [6, Proposition 4.16]):
@@
3.  The Geometric Integration by Parts Formula
@@
Let $(M,g)$ be a Riemannian manifold with boundary $\partial M$, and let $\nabla$ denote the Levi-Civita connection on $TM$. Let $E\to M$ be a smooth vector bundle equipped with a fiber metric $h_{E}$ and a connection $\nabla^{E}$ compatible with $h_{E}$. We continue to denote by $\nabla$ the induced connection on the tensor bundles $T^{(k,l)}(TM)\otimes E$.
@@
For $\mathrm{(II)}$, the term inside $\partial_{I_{s}}$ contains $G$ algebraically (zero derivatives). Thus, applying $\partial_{I_{s}}$ yields terms with at most $s$ derivatives on $G$. Similarly for $\mathrm{(III)}$, the dependence on $G$ is linear without additional derivatives, so expanding yields terms of order at most $s$. Combining these, there exist coefficients $\bigl(\widetilde{B}_{\beta}\bigr)$ such that:
@@
4.  Local-to-Global Estimates and Structural Theorems
Having established the global integration by parts formula and the local structure of the formal adjoint, we now turn to the fundamental structural theorems of the theory. Our primary objective in this section is to establish a characterization of Sobolev sections on a coordinate chart in terms of their component functions with respect to a local frame (Lemma 3); this allows us to reduce intrinsic analytic problems to their Euclidean counterparts. To achieve this, we require a result ensuring norm equivalence for fiber metrics on vector bundles over compact manifolds. A similar equivalence is well known for the tangent bundle (see, e.g., [6]). This behavior extends to the Sobolev setting, where different choices of metric and connection yield equivalent Sobolev norms on compact manifolds, as we have previously established in [9, Theorem 2] for Sobolev spaces defined via completions.
@@ 4.  Local-to-Global Estimates and Structural Theorems
Let $M$ be a smooth compact manifold (with or without boundary) and let $E\to M$ be a smooth vector bundle of finite rank. Let $h$ and $\widetilde{h}$ be two smooth fiber metrics on $E$. Then, there exist constants $c,C>0$ such that for all $p\in M$ and all $v\in E_{p}$,
@@
In local coordinates, writing $F=F^{a}e_{a}$ and $\Psi=\Psi^{c}_{J_{s}}\,dx^{J_{s}}\otimes e_{c}$, this identity reads
@@
This expression matches the expansion of $\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle\,d\lambda_{g}$ exactly. Therefore, $G$ is the $s$-th weak covariant derivative of $F$ in $U$. Moreover, $G\in L^{p}(T^{(0,s)}(TU)\otimes E|_{U})$ because each coordinate component is a linear combination of weak derivatives of the components of $F$ (up to order $s$) multiplied by smooth bounded functions. Since $s\leq m$ was arbitrary, $F|_{U}\in W^{m,p}\big(E|_{U}\big)$.
@@
Define a fiber metric $\overline{h}_{E}$ on $E|_{U}$ by $\overline{h}_{E}(u,v)=\displaystyle\sum_{a=1}^{r}u^{a}v^{a}$ for sections $u=\displaystyle\sum_{a=1}^{r}u^{a}e_{a},v=\sum_{a=1}^{r}v^{a}e_{a}$. Also consider the Riemannian metric $\delta:=\phi^{-1}\overline{g}$ on $\overline{U}$. Since $\overline{U}$ is compact, Lemma 2 provides constants $c_{0},C_{0}>0$ for the pointwise equivalence of the fiber norms, uniform for all $s\in\{0,\dots,m\}$. By integrating these estimates and observing that the Riemann–Lebesgue measures $d\lambda_{g}$ and $d\lambda_{\overline{g}}$ are comparable on compact sets through positive constants [9, Lemma 2], we conclude that there exist constants $c,C>0$ such that the following equivalence of Sobolev norms holds:
@@
5.  Global Approximation Results
@@
By Theorem 2, $W^{m,p}(E)$ coincides with the completion-based space $H^{m,p}(E)$. The independence of $H^{m,p}(E)$ from the underlying geometric data on compact manifolds follows by generalizing the analytic arguments established for the scalar-valued case in [9, Theorem 2] to the vector bundle setting.∎
6.  Sobolev Embeddings and the Rellich–Kondrashov Theorem

### v2

@@ Abstract
This article develops a unified and intrinsic framework for the theory of Sobolev spaces on vector bundles over Riemannian manifolds. The analytical core of our approach is an explicit higher-order geometric integration by parts formula, which characterizes the formal adjoint of the covariant derivative as a global differential operator. This identity is established on arbitrary Riemannian manifolds with boundary, without assuming completeness or compactness.
While first-order integration by parts identities are classical, explicit higher-order formulas with precise boundary terms are rarely stated in the literature. As applications of this framework, we recover the classical Meyers–Serrin theorem on arbitrary manifolds and, in the compact case, the Sobolev embedding and Rellich–Kondrachov compactness theorems, providing direct and self-contained proofs. At the end of this work we also stablish a Green Formula and we use it to stablish norm equivalence in Sobolev Spaces on vector bundles with closed manifold as base space and the Bochner laplacian operator. As a corollary we show that, in case of trivial vector bundles this equivalence reduces to a well known (but non proved rigorously in literature) result for closed manifolds and the Laplace-Beltrami operator.
By emphasizing intrinsic global arguments and sharp local-to-global norm equivalence estimates, rather than ad hoc coordinate patching, this work offers a transparent and accessible foundation for the study of Sobolev spaces on vector bundles, suitable for researchers in global analysis, differential geometry, and partial differential equations.
keywords:
@@
Sobolev spaces, Vector bundles, Geometric integration by parts, Embedding theorems, Rellich-Kondrashov theorem
pacs:
MSC Classification]46E35, 58A99, 58C99
1 Introduction
@@
The analysis of Sobolev spaces on vector bundles is a fundamental pillar of modern geometric analysis. A central issue in this theory is ensuring the equivalence between the two standard definitions of these spaces: the completion of the space of smooth sections with respect to the Sobolev norm, denoted by $H^{m,p}(E)$, and the space of sections possessing weak covariant derivatives in $L^{p}$, denoted by $W^{m,p}(E)$.
On Riemannian manifolds, the treatment of this equivalence and the associated structural results of Sobolev theory vary significantly across the literature. Standard references such as [1] and [2] often introduce Sobolev spaces as completions of smooth sections, while more intrinsically oriented approaches, such as [3], define them via distributional covariant derivatives. Establishing the identity $H^{m,p}(E)=W^{m,p}(E)$—the Meyers–Serrin theorem, which is classical in the scalar case—is therefore essential to reconcile these perspectives. For general vector bundles, this equivalence is also known and has been established in a very general setting by Guidetti [4] using the theory of linear partial differential operators and the theory of elliptic operators. We will also derive this result in an alternative manner, using only a geometrical intrinsic approach.
Beyond density, the fundamental theorems of Sobolev theory—embedding and compactness—are likewise well established in the literature for vector bundle–valued sections, particularly in the compact case. The work of Palais [5] remains a foundational reference for Sobolev embeddings and Rellich–Kondrachov compactness on compact manifolds. However, existing proofs are often scattered across different sources, rely on diverse analytical frameworks, or involve constructions that can obscure the underlying geometric structure, making them less accessible to a broader audience interested in both partial differential equations and differential geometry.
The primary purpose of this paper is not to reestablish these classical results, but rather to provide a global and intrinsic integration by parts formula for higher-order covariant derivatives acting on sections of smooth vector bundles over Riemannian manifolds with boundary (Theorem 2). This formula constitutes the analytical core of the work. Its proof is carried out in a fully geometric manner and does not rely on elliptic regularity, completeness, or compactness assumptions. By formulating the identity intrinsically and making all boundary terms explicit, we obtain a transparent characterization of the formal adjoint of the covariant derivative as a global differential operator.
As applications of this integration by parts formula, we revisit several classical results in the theory of Sobolev spaces on vector bundles. In Section 5, we recover the Meyers–Serrin theorem on general Riemannian manifolds by a direct argument based on the integration by parts identity and the classical scalar Meyers–Serrin theorem. The resulting statement for vector bundle–valued sections is in full agreement with the corresponding extensions available in the literature.
In the compact setting, treated in Section 6, we apply the same framework to provide streamlined proofs of the classical Sobolev embedding theorem and the Rellich–Kondrachov compactness theorem, fully consistent with known results.
By placing the integration by parts formula at the center of the theory, this work offers a unifying and geometrically transparent perspective on Sobolev spaces on vector bundles. The approach emphasizes global and intrinsic arguments and is designed to be accessible to a broad audience, including researchers in global analysis, differential geometry, and partial differential equations. We expect the framework developed here to serve both as a conceptual reference and as a flexible tool for further investigations involving vector bundles and boundary value problems.
2 Notation and Preliminaries
@@ 2.2 Sobolev Spaces
We begin with the precise definitions of the Sobolev spaces used in this work.
@@
Recall that for the tensor field $F^{(a)}$, the connection components are given by (see, e.g., [7, Proposition 4.16]):
@@
3 The Geometric Integration by Parts Formula
@@
Let $(M,g)$ be a Riemannian manifold with boundary $\partial M$, and let $\nabla$ denote the Levi-Civita connection on $TM$. Let $E\to M$ be a smooth vector bundle equipped with a fiber metric $h_{E}$ and a connection $\nabla^{E}$ that is compatible with $h_{E}$. We continue to denote by $\nabla$ the induced connection on the tensor bundles $T^{(k,l)}(TM)\otimes E$.
@@
For $\mathrm{(II)}$, the term inside $\partial_{I_{s}}$ contains $G$ algebraically (zero derivatives). Thus, applying $\partial_{I_{s}}$ yields terms with at most $s$ derivatives on $G$. Similarly, for $\mathrm{(III)}$, the dependence on $G$ is linear without additional derivatives, so expanding yields terms of order at most $s$. Combining these, there exist coefficients $\bigl(\widetilde{B}_{\beta}\bigr)$ such that:
@@
4 Local-to-Global Estimates and Structural Theorems
Having established the global integration by parts formula and the local structure of the formal adjoint, we now turn to the fundamental structural theorems of the theory. Our primary objective in this section is to establish a characterization of Sobolev sections on a coordinate chart in terms of their component functions with respect to a local frame (Lemma 5); this allows us to reduce intrinsic analytic problems to their Euclidean counterparts. To achieve this, we require a result ensuring norm equivalence for fiber metrics on vector bundles over compact manifolds. A similar equivalence is well known for the tangent bundle (see, e.g., [7]). This behavior extends to the Sobolev setting, where different choices of metric and connection yield equivalent Sobolev norms on compact manifolds, as we have previously established in [8, Theorem 2] for Sobolev spaces defined via completions.
@@ 4 Local-to-Global Estimates and Structural Theorems
Let $M$ be a smooth compact manifold (with or without boundary), and let $E\to M$ be a smooth vector bundle of finite rank. Let $h$ and $\widetilde{h}$ be two smooth fiber metrics on $E$. Then, there exist constants $c,C>0$ such that for all $p\in M$ and all $v\in E_{p}$,
@@
In local coordinates, writing $F=F^{a}e_{a}$ and $\Psi=\Psi^{c}_{J_{s}}\,dx^{J_{s}}\otimes e_{c}$, this identity reads:
@@
This expression matches the expansion of $\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle\,d\lambda_{g}$ exactly. Therefore, $G$ is the $s$-th weak covariant derivative of $F$ in $U$. Moreover, $G\in L^{p}(T^{(0,s)}(TU)\otimes E|_{U})$ because each coordinate component is a linear combination of weak derivatives of the components of $F$ (up to order $s$) multiplied by smooth, bounded functions. Since $s\leq m$ was arbitrary, $F|_{U}\in W^{m,p}\big(E|_{U}\big)$.
@@
Define a fiber metric $\overline{h}_{E}$ on $E|_{U}$ by $\overline{h}_{E}(u,v)=\displaystyle\sum_{a=1}^{r}u^{a}v^{a}$ for sections $u=\displaystyle\sum_{a=1}^{r}u^{a}e_{a},v=\sum_{a=1}^{r}v^{a}e_{a}$. Also consider the Riemannian metric $\delta:=\phi^{-1}\overline{g}$ on $\overline{U}$. Since $\overline{U}$ is compact, Lemma 4 provides constants $c_{0},C_{0}>0$ for the pointwise equivalence of the fiber norms, uniform for all $s\in\{0,\dots,m\}$. By integrating these estimates and observing that the Riemann–Lebesgue measures $d\lambda_{g}$ and $d\lambda_{\overline{g}}$ are comparable on compact sets through positive constants [8, Lemma 2], we conclude that there exist constants $c,C>0$ such that the following equivalence of Sobolev norms holds:
@@
5 Global Approximation Results
@@
By Theorem 8, $W^{m,p}(E)$ coincides with the completion-based space $H^{m,p}(E)$. The independence of $H^{m,p}(E)$ from the underlying geometric data on compact manifolds follows by generalizing the analytic arguments established for the scalar-valued case in [8, Theorem 2] to the vector bundle setting.∎
6 Sobolev Embeddings and the Rellich–Kondrashov Theorem